# Multi-SpaCE: Multi-Objective Subsequence-based Sparse Counterfactual Explanations for Multivariate Time Series Classification

### v1

@@ Appendix C Sparsity and Contiguity results
As discussed in Section 4.5, sparsity and contiguity must be jointly evaluated. Sparsity minimizes the number of changes applied to the original instance, while contiguity ensures that these changes are coherent. Several methods impose constraints on the structure of changes. For example, NG and COMTE restrict modifications to a single subsequence per channel, resulting in counterfactuals with a low number of subsequences. This is evident in Table 16, where these methods achieve the lowest Number of Subsequences scores. Glacier also produces counterfactuals with a single subsequence in the univariate setting, although this outcome is due to altering every point of the original instance, leading to the worst sparsity scores overall. In the multivariate setting, Multi-SpaCE achieves the second lowest rank, while on the univariate setting, it achieves the third lowest rank, tying with AB-CF.
(a) NoS for Multivariate Datasets
@@
AWR 3.07 17.53 31.58 18.07
BasicMotions 2.9 12.33 25.92 6.9
Cricket 3.25 9.55 - 41.69
Epilepsy 4.55 9.48 27.43 6.01
NATOPS 3.27 53.02 35.86 10.34
PEMS-SF 52.03 1943.78 - 1123.77
PenDigits 2.68 3.24 1.29 1.71
RacketSports 2.92 13.16 13.68 2.03
SR-SCP1 5.26 15.56 226.06 47.22
UWave 2.64 6.33 29.49 3.12
Average Rank 1.3 2.9 3.5 2.2
@@
Coffee 1.0 1.0 1.0 1.82 10.77 2.86
ECG200 1.0 1.0 1.0 1.68 6.02 2.04
FordA 1.0 1.0 1.0 2.21 16.35 3.54
Gunpoint 1.0 1.0 1.0 1.65 5.87 1.51
HandOutlines 1.0 1.0 1.0 2.56 113.63 6.38
ItalyPower 1.0 1.0 1.0 2.39 1.6 1.17
PPOC 1.0 1.0 1.0 1.62 4.49 1.44
Strawberry 1.0 1.0 1.0 2.12 11.83 1.83
CBF 1.0 - - 2.11 6.7 1.64
CinCECGTorso 1.0 - - 1.75 40.52 2.11
TwoPatterns 1.0 - - 2.23 4.78 1.16
ECG5000 1.0 - - 1.25 4.94 2.19
Plane 1.0 - - 2.61 5.06 2.64
FacesUCR 1.0 - - 2.49 4.92 1.98
NI-ECG2 1.0 - - 1.58 19.25 4.09
Average Rank 1.0 1.0 1.0 3.6 5.0 3.6
@@
Imposing restrictions on the form of changes can result in requiring more alterations to achieve the desired outcome, negatively affecting sparsity. Consequently, NG and COMTE perform worse than Multi-SpaCE in terms of sparsity, as shown in Table 17. Multi-SpaCE consistently achieves the best sparsity performance across univariate datasets and has the lowest rank for multivariate datasets.
@@
AWR 0.34 0.47 0.09 0.18
BasicMotions 0.48 0.55 0.13 0.19
Cricket 0.5 0.48 - 0.22
Epilepsy 0.77 0.46 0.24 0.16
NATOPS 0.14 0.75 0.05 0.11
PEMS-SF 0.04 0.54 - 0.11
PenDigits 0.75 0.65 0.36 0.28
RacketSports 0.48 0.59 0.15 0.07
SR-SCP1 0.5 0.61 0.12 0.11
UWave 0.88 0.43 0.23 0.12
Average Rank 3.1 3.4 1.62 1.4
@@
Coffee 0.56 1.0 1.0 0.43 0.47 0.08
ECG200 0.72 1.0 1.0 0.44 0.46 0.12
FordA 0.46 1.0 1.0 0.58 0.52 0.06
Gunpoint 0.42 1.0 1.0 0.35 0.37 0.11
HandOutlines 0.47 1.0 1.0 0.69 0.47 0.06
ItalyPower 0.61 1.0 1.0 0.56 0.46 0.15
PPOC 0.61 1.0 1.0 0.86 0.38 0.07
Strawberry 0.38 1.0 1.0 0.69 0.45 0.04
CBF 0.36 - - 0.39 0.42 0.13
CinCECGTorso 0.15 - - 0.54 0.38 0.06
TwoPatterns 0.38 - - 0.51 0.34 0.1
ECG5000 0.83 - - 0.71 0.5 0.31
Plane 0.51 - - 0.5 0.61 0.18
FacesUCR 0.38 - - 0.43 0.56 0.14
NI-ECG2 0.44 - - 0.66 0.54 0.05
Average Rank 2.87 5.0 5.0 3.2 2.87 1.0
@@
This relation between contiguity and sparsity motivated the need for their joint evaluation, as highlighted in Section 4.5. Multi-SpaCE demonstrates clear superiority over other methods in the literature when both metrics are considered together.

### v2

@@ Appendix C Sparsity and Contiguity results
As discussed on the main paper, sparsity and contiguity must be jointly evaluated. Sparsity minimizes the number of changes applied to the original instance, while contiguity ensures that these changes are coherent. Several methods impose constraints on the structure of changes. For example, NG and COMTE restrict modifications to a single subsequence per channel, resulting in counterfactuals with a low number of subsequences. This is evident in Table 12, where these methods achieve the lowest Number of Subsequences scores. Glacier also produces counterfactuals with a single subsequence in the univariate setting, although this outcome is due to the change of every point in the original instance, leading to the worst sparsity scores overall. In the multivariate setting, Multi-SpaCE achieves the second lowest rank. Regarding the univariate setting, it achieves the fourth lowest rank, ranking after NG Glacier(AE) and Glacier methods which use a single subsequence of changes.
(a) Number of Subsequences (NoS) for Multivariate Datasets
@@
AWR 2.97 10.6 52.78 7.03
BasicMotions 3.0 24.11 15.33 4.4
Cricket 3.0 6.88 - 57.62
Epilepsy 4.47 7.87 21.95 5.56
NATOPS 3.24 47.28 33.18 10.44
PEMS-SF 86.81 1241.13 - 1230.05
PenDigits 2.87 3.13 1.31 1.68
RacketSports 2.78 8.23 10.88 2.97
SR-SCP1 3.93 10.69 354.09 19.39
UWave 2.58 6.01 30.5 3.69
Average Rank 1.2 3.1 3.38 2.2
@@
Coffee 1.0 1.0 1.0 1.82 10.93 2.96
ECG200 1.0 1.0 1.0 1.99 5.71 1.39
FordA 1.0 1.0 1.0 2.39 10.85 4.15
Gunpoint 1.0 1.0 1.0 1.91 4.95 1.48
HandOutlines 1.0 1.0 1.0 2.37 90.98 2.48
ItalyPower 1.0 1.0 1.0 2.67 1.79 1.22
PPOC 1.0 1.0 1.0 1.78 5.26 1.35
Strawberry 1.0 1.0 1.0 1.82 10.81 1.86
CBF 1.0 - - 2.24 5.75 1.75
CinCECGTorso 1.0 - - 2.26 25.7 3.9
TwoPatterns 1.0 - - 2.15 5.12 1.4
ECG5000 1.0 - - 1.57 5.41 1.2
Plane 1.0 - - 1.8 5.74 1.9
FacesUCR 1.0 - - 2.9 3.96 2.21
NI-ECG2 1.0 - - 1.99 13.39 3.5
Average Rank 1.0 1.0 1.0 3.67 5.0 3.53
@@
Imposing restrictions on how to change the original instance to form the counterfactual can result in requiring more alterations to achieve the desired outcome, negatively affecting sparsity. Consequently, NG and COMTE perform worse than Multi-SpaCE in terms of sparsity, as shown in Table 13. Multi-SpaCE consistently achieves the best sparsity performance across univariate datasets and share the lowest average rank with DiscoX for multivariate datasets.
@@
AWR 0.33 0.9 0.11 0.19
BasicMotions 0.5 0.56 0.12 0.2
Cricket 0.5 0.97 - 0.39
Epilepsy 0.8 0.77 0.27 0.32
NATOPS 0.14 0.73 0.04 0.12
PEMS-SF 0.06 0.49 - 0.15
PenDigits 0.82 0.72 0.35 0.29
RacketSports 0.46 0.93 0.13 0.12
SR-SCP1 0.33 0.87 0.14 0.06
UWave 0.86 0.53 0.27 0.17
Average Rank 3.0 3.5 1.5 1.5
@@
Coffee 0.22 1.0 1.0 0.51 0.3 0.09
ECG200 0.34 1.0 1.0 0.48 0.46 0.11
FordA 0.43 1.0 1.0 0.64 0.58 0.09
Gunpoint 0.23 1.0 1.0 0.56 0.37 0.14
HandOutlines 0.31 1.0 1.0 0.48 0.39 0.01
ItalyPower 0.71 1.0 1.0 0.55 0.51 0.26
PPOC 0.45 1.0 1.0 0.53 0.35 0.06
Strawberry 0.41 1.0 1.0 0.55 0.42 0.06
CBF 0.31 - - 0.44 0.36 0.2
CinCECGTorso 0.27 - - 0.5 0.37 0.19
TwoPatterns 0.29 - - 0.59 0.3 0.11
ECG5000 0.26 - - 0.81 0.52 0.19
Plane 0.42 - - 0.57 0.55 0.21
FacesUCR 0.37 - - 0.57 0.61 0.19
NI-ECG2 0.33 - - 0.52 0.54 0.09
Average Rank 2.2 5.0 5.0 3.8 3.0 1.0
@@
This relation between contiguity and sparsity motivated the need for their joint evaluation. Multi-SpaCE demonstrates superiority over other methods in the literature when both metrics are considered together.